OpenStudy (hybrik):

Simplify: log2^25 + log 4^4

OpenStudy (hybrik):

log 2^25 + log 4^5 my bad

OpenStudy (jdoe0001):

\(\large log_225+log_45?\)

jhonyy9 (jhonyy9):

do you know the property of logarithm ?

jhonyy9 (jhonyy9):

log a^x = xlog a for example

jhonyy9 (jhonyy9):

or log a + log b = log (a*b) do you know these ?

jhonyy9 (jhonyy9):

without them you not can solving exercises like this never

jhonyy9 (jhonyy9):

@hybrik what is your opinion please ?

OpenStudy (kayders1997):

Hi

OpenStudy (hybrik):

@jhonyy9 Sorry I had to shower, Im good at math but my weakness is logarithims, I do know it

jhonyy9 (jhonyy9):

ok. so how can i help you ?

OpenStudy (hybrik):

just dont know how to simplify

jhonyy9 (jhonyy9):

check please the above wrote property and try getting from there what formula you can us here ? wait

jhonyy9 (jhonyy9):

but sorry first in the left term there is 2 like base and in the right term there is 4 like base term so how can you make the 4 to be same with 2 using exponent ?

jhonyy9 (jhonyy9):

log 2^25 + log 4^5 = 4= 2^2 yes ?

OpenStudy (hybrik):

yea

jhonyy9 (jhonyy9):

so 4^5 = (2^2)^5 = ?

OpenStudy (hybrik):

yea

jhonyy9 (jhonyy9):

= 2^?

OpenStudy (hybrik):

2log2^5?

jhonyy9 (jhonyy9):

first without log please answer this question

OpenStudy (hybrik):

= 2^? one?

jhonyy9 (jhonyy9):

yes what number will get in place of ?

OpenStudy (hybrik):

idk what your asking

OpenStudy (hybrik):

2^? = 2^x?

jhonyy9 (jhonyy9):

(x^a)^b = x^(?)

jhonyy9 (jhonyy9):

property of exponents

OpenStudy (hybrik):

idk

OpenStudy (hybrik):

a x^b?

jhonyy9 (jhonyy9):

why ? you have said me that you dont know the logarithm and now you say that you dont know the property of exponents - so when wan you learning them ?

OpenStudy (hybrik):

idk im in 6th grade

OpenStudy (hybrik):

i teach myself, logx2b = xlogxb?

jhonyy9 (jhonyy9):

but you have learned about exponents - yes ?

OpenStudy (hybrik):

yea

jhonyy9 (jhonyy9):

sorry but this logx2b i dont understand how you have wrote there

OpenStudy (hybrik):

just continue ur x2ab thing

jhonyy9 (jhonyy9):

so when there are exponent on exponent so the exponents are coming multiplied (x^a)^b = x^(a*b) ok. ? remember and learn it please

OpenStudy (hybrik):

oh yea i know that

OpenStudy (hybrik):

misread it my bad

jhonyy9 (jhonyy9):

now this is easy to say that i know sorry

jhonyy9 (jhonyy9):

so how you can continue it your exercise ?

OpenStudy (hybrik):

idk

jhonyy9 (jhonyy9):

what you have got ? in the last step ?

OpenStudy (hybrik):

2log2^5

OpenStudy (hybrik):

cause log4 ^ 5 = log 2^2 ^5

OpenStudy (hybrik):

and you can move the 2 to have 2log2 ^ 5

jhonyy9 (jhonyy9):

log 2^25 + log 4^5 = no first i have said that without logarithm 4^5 = 2^(?)

OpenStudy (hybrik):

4^5 = 2^2^5

OpenStudy (hybrik):

(2^2)^5

jhonyy9 (jhonyy9):

2^(?)

jhonyy9 (jhonyy9):

because in the first term there is 2 on exponent so you need using in the second 2 on exponent - do you understand it now why ?

jhonyy9 (jhonyy9):

2^2^5 = 2^(2*5) =?

OpenStudy (hybrik):

yea

OpenStudy (hybrik):

wait no

OpenStudy (hybrik):

wait yes

jhonyy9 (jhonyy9):

please write what i ask you

jhonyy9 (jhonyy9):

2^(2*5) = ?

OpenStudy (hybrik):

oh i didnt see that message sorry

OpenStudy (hybrik):

1024

jhonyy9 (jhonyy9):

nonono

jhonyy9 (jhonyy9):

the base need being 2

OpenStudy (hybrik):

it is already 2^10?

jhonyy9 (jhonyy9):

exactly

jhonyy9 (jhonyy9):

log 2^25 +log 2^10 = log ? than you know this formula log a +log b = log a*b

OpenStudy (hybrik):

yeah

OpenStudy (hybrik):

uhm = log 2 250

OpenStudy (hybrik):

where log 2 ^10 come from

OpenStudy (hybrik):

did u move the 2 on the 5, wouldnt you move it to be 2log2^5 and not log2^10?

jhonyy9 (jhonyy9):

check above 4^5 = 2^2^5 = 2^(2*5) = 2^10 yes ?

OpenStudy (hybrik):

yes

jhonyy9 (jhonyy9):

because in the first term there is 2^25 - yes ?

jhonyy9 (jhonyy9):

so please log 2^25 +log 2^10 = log ? use the above wrote formula please

OpenStudy (hybrik):

wait r u simplifying log 4^5 or log 2^25?

OpenStudy (hybrik):

@jhonyy9 i said log 2 250

jhonyy9 (jhonyy9):

why ? how you know a^x *a^y = ?

OpenStudy (hybrik):

u multiply the exponents when adding, so log 25*10 = log 2 250?

jhonyy9 (jhonyy9):

no multiplie the exponents just when there are exponents on exponents

jhonyy9 (jhonyy9):

in this case of multiplie the same terms with exponents just assum the exponents property of exponents - sorry what you need learning again i see

jhonyy9 (jhonyy9):

so how will be ?

jhonyy9 (jhonyy9):

2^25 *2^10 = 2^?

jhonyy9 (jhonyy9):

please cooperate because i need to go

jhonyy9 (jhonyy9):

where are you from please ? because here in Europe is in the night 3,12

jhonyy9 (jhonyy9):

are you here ?

OpenStudy (hybrik):

was eating sorry?

OpenStudy (hybrik):

@jhonyy9 2^35

jhonyy9 (jhonyy9):

ok

jhonyy9 (jhonyy9):

so now there you have got log 2^35 = ? using this formula log a^x = xlog a

OpenStudy (hybrik):

35 log2?

jhonyy9 (jhonyy9):

35log 2 hence please because than you write 35 log2 so this mean 35 space log base 2 of nothing

OpenStudy (hybrik):

ok, its not an answer choice so its wrong?

OpenStudy (hybrik):

r u sure u can just do 2log2^5 + log2^25

jhonyy9 (jhonyy9):

what are the choices ?

OpenStudy (hybrik):

\[2/5 \log_{5}2 , 2/5 \log_{2}5, 5/2\log_{5}2, 5/2\log_{2}5 \]

jhonyy9 (jhonyy9):

ok but till this moment you dont used nothing base of logarithm - why ?

OpenStudy (hybrik):

did u read my question wrong

OpenStudy (hybrik):

\[\log_{2}25 + \log_{4}5\]

jhonyy9 (jhonyy9):

than you check in the first answer @jdoe have asked you how is this exercise correct ?

OpenStudy (hybrik):

thats my question